OpenStudy (anonymous):

Reposting this because I think I asked in a confusing manner before. Im using sum & Difference identities will post more in the response because I cant figure out how to put the proper symbols in here lmfao.

OpenStudy (anonymous):

Problem= \[\cos (11\pi \div 12)\] should I use \[\cos (\alpha-\beta) = \cos(\alpha)\cos (\beta) +\sin (\alpha) \sin(\beta) \] Im assuming I am doing this right and (\alpha) is 45 and (\beta) is -30 ??